# Policymaker meetings as heteroscedasticity shifters:

### v1

@@ A.2 Smoothed LP
Figure 10: Identification of monetary policy shock in the US. Smoothed IRF estimates with point-wise and simultaneous bands. Heteroscedasticity-IV smoothed LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dot-dash line)
@@ A.3 Bonferroni bounds
Figure 11: Identification of monetary policy shock in the US. Sup-t vs Bonferroni simultaneous bands. Heteroscedasticity-IV LP with $k=0,p=12$ estimator (blue solid line), Bonferroni bounds (gold dots) and simultaneous sup-t bounds (dot-dash line)
@@ A.4 Comparison with conventional VAR approch
Figure 12: Identification of monetary policy shock in the US. LP vs VAR using Christiano et al. (1999). Cholesky identification with LP estimator with $k=0,p=12$ (blue line) and homoscedastic VAR estimator (violet solid) with corresponding 68% point-wise bands ( gold dashes and violet dots)
@@ B.1 Proof of Theorem 1
$b_{\cdot 1}^{\prime}\Sigma^{-1}_{\eta}b_{\cdot 1}=1$ (48)
Step 2. The reduced form shocks have the following representation for any $k=1,\dots,n$
$\eta_{1t}\eta_{kt}$ $=b_{k1}b_{11}\varepsilon^{2}_{1t}+\sum_{i\neq j}b_{ki}b_{1j}\varepsilon_{it}\varepsilon_{jt}+\sum_{i\neq 1}b_{ki}b_{1i}\varepsilon^{2}_{it}$ (49)
@@
$=b_{\cdot 1}b_{11}\mathbb{E}(\varepsilon^{2}_{1t}-1)(Z_{t}-\mathbb{E}Z_{t})=b_{\cdot 1}b_{11}\rho$ (52)
@@
$\sqrt{\gamma^{\prime}\Sigma^{-1}_{\eta}\gamma}=|b_{11}\rho|$ (53)
@@
$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma\operatorname{sign}(b_{11}\rho).$ (54)
@@ B.2 Proof of LP consistency
Suppose that Assumptions 2 and 3 hold. Consider LS residuals $\hat{\eta}_{t}(H)$ from $H-$step forward autorgression,
@@
The proof proceeds in 3 steps. First, we show that (12) can be decomposed into linearly independent univariate AR processes using Jordan decomposition. Then, we show that the LS estimators for individual processes have well-defined asymptotic distributions. Finally, we use invariance of LS estiates with respect to linear transformations to estimate the residual in (56).
Step 1. Notice that $\mu_{t}=(1,t,\dots,t^{k})^{\prime}$ satisfies a first order deterministic difference equation
$\mu_{t}=Q_{\mu}\mu_{t-1}$ (57)
@@
Matrix $Q_{y}$ can be represented in Jordan normal form  [^18] See, for example, Gantmakher (2000).
$Q_{y}=P_{y}J_{y}P_{y}^{-1}.$ (63)
Then, following ideas in Tsay and Tiao (1990) and Sims et al. (1990), we get VAR(1) representation
@@
with $\lambda_{i}$ being $i$-th eigenvalue of $Q_{y}$ (not necessarily distinct) with geometric multiplicity $n_{i}$. Suppose, without loss of generality, that the blocks $J_{y,i}$ are arranged in decreasing order according to the absolute values of the corresponding $\lambda_{i}$. If unit root is present, then start with $\lambda_{1}=1$. By Assumption 3, there are no roots larger than 1 in absolute value.
@@
$\tilde{X}_{t}\triangleq\begin{pmatrix}P_{x}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}\triangleq F(L)\nu_{t}$ (66)
@@
and $\nu_{t}=(\eta^{\prime}_{t},(\xi^{1}_{t}(\lambda_{1}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{1}))^{\prime},(\xi^{1}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{1}_{t}(\lambda_{b}))^{\prime},\dots,(\xi^{g_{b}}_{t}(\lambda_{1}))^{\prime},1,t,\dots,t^{k})^{\prime}$ with $\xi^{j}_{t}(\lambda)\triangleq\lambda\sum_{s=1}^{t}\xi^{j-1}_{t}(\lambda)$, $\xi^{0}_{t}(\lambda)\triangleq\eta_{t}\lambda^{t}$, $i=1,\dots,b$ $\lambda_{i}$ are unit roots of (12) and $g_{i}$ are their corresponding maximum geometric multiplicities. The blocks $F_{i,j}$ corresponding to two different unit roots $\lambda_{i}$ and $\lambda_{j}$ are 0. The lag matrix $F_{11}(L)$ is obtained by inversion of the stationary rows of (64) and therefore $\sum^{\infty}_{j=0}j^{g}|F_{11j}|<\infty$ and $\sum_{j=0}^{\infty}F_{11j}F_{11j}^{\prime}$ is not singular.
Following Sims et al. (1990); Tsay and Tiao (1990), one can arrange the properly chosen powers of $T^{1/2}$ in a diagonal matrix $\Upsilon_{T}$ such that $\sum_{t=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{t-1}$ have a bounded variance (for example, for stationary components it is $T^{1/2}$, for components containing $\xi^{j}_{t}(\lambda_{2}))$ it is $T^{(1+j)/2}$ etc.).
@@
Consider $h=0$ first. The proof for deterministic, stationary and ordinary unit root components follows the proofs of Lemmas 1 and 2 in Sims et al. (1990) with a modification required to account for conditionally heteroscedastic shocks $\eta_{t}$. Namely, proof of Lemma 1 Sims et al. (1990) requires replacement of Functional CLT for sums of $\xi^{1}_{t}$ in part (a) and Theorem 2.4(ii) from Chan and Wei (1988) in part (c) for case $p=1$. Under Assumption 2, Theorem 3.1 (De Jong and Davidson, 2000) applied to $\xi^{1}_{t}$ gives the required FCLT under mixing processes assumption. (De Jong and Davidson, 2000, Theorem 4.1) for any pair of components of $\xi^{1}_{t}$ replaces Theorem 2.4(ii) of Chan and Wei (1988) for mixing case (white noise assumption on $\eta_{t}$ ensures that we do not need recenter the sums). For complex unit roots, the limiting distribution is given in Theorem 4.4 of Tsay and Tiao (1990), which also requires replacing FCLT for martingale differences with mixing white noise (De Jong and Davidson, 2000, Theorem 4.1) and stong law of large numbers used in from Chan and Wei (1988) for MDS with Corollary 3.2 in Rio et al. (2017). For our purposes, we do not need explicit form of $\phi$ and $\tilde{V}$.
@@
where $\Theta(h)\triangleq(A_{1}(h),A_{2}(h),\dots,A_{p}(h),V(h))$ and $X_{t-1}\triangleq vec(y_{t-1},y_{t-2},\dots,y_{t-p},1,(t-1),\dots(t-1)^{k})$.
@@
$\hat{\eta}_{t+h}(h)-\eta_{t+h}(h)=(\Theta(h)-\hat{\Theta}(h))X_{t-1}=-X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h).$ (71)
@@
$X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h)$ $=\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime}(P^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime})^{-1}\sum_{r=p}^{T-h}(P^{-1})\tilde{X}_{r-1}\eta_{r+h}(h),$ (72)
$=\tilde{X}_{t-1}^{\prime}(\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{t-1}^{\prime})^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\eta_{r+h}(h).$ (73)
@@
$=C_{h}^{\prime}+\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\frac{1}{T}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))+O_{p}(T^{-1})$ (82)
Since $\eta_{t}$ is stationary and mixing and thus sequence $\eta_{t}\eta_{t+h}$ is ergodic with finite fourth moments, so we can apply the ergodic theorem to get the desired result.
@@ B.3 Proof of consistency of IV estimtor
$\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})\eta_{jt}=O_{p}(1)$ (85)
Next, by Assumption 2.5, $Z_{t}\eta_{t}$ is zero meand and is uncorrelated with $\tilde{X}_{t-1}^{\prime}$. Using the same argument as in Step 2 of the proof of Lemma B.1, $\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}$ is a tight random sequence by De Jong and Davidson (2000, Theorem 4.1). So
$(\sum_{t=p}^{T}(\hat{\eta}_{t}-\eta_{t})\eta^{\prime}_{t}Z_{t})^{\prime}=\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r}=O_{p}(1)$ (86)
The last term in (84) can be bounded by Cauchy-Schwarz theorem,
@@
$\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\bar{Z}_{T})$ $=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(Z_{t}-\mathbb{E}Z_{t})\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})$ (92)
$=\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(1)$ (93)
@@ B.4 Proof of asymptotic normality of the LP estimators
By Assumption 2, $\Xi^{H}_{t}$ is a strong mixing process of size at least $-2(1+\epsilon)/\epsilon$ with zero mean and finite $2+\epsilon$ moments (condition DMR is satisfied, see p. 12 in Rio et al. (2017)). By Theorem 4.2 in Rio et al. (2017) and the Cramer-Wold theorem, there exist the long run variance matrix $\Omega$ for $\Xi^{H}_{t}$ and
@@ B.5 Proof of consistency of HAC standard errors
Step 1.To show consitency of the whole matrix $\hat{\Omega}=\hat{\Omega}(H_{1})$, we need to establish consistency of HAC estimators for all combinations $a^{\prime}\Xi_{t}$. Then the HAC estimator for particular $a$ is defined as
$\hat{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\hat{\Xi}_{t})(\hat{\Xi}_{t+\ell}^{\prime}a)$ (99)
@@
$\tilde{\Omega}(a)=\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\Xi_{t})(\Xi_{t+\ell}^{\prime}a)$ (100)
@@
Step 2. What remains to show is $\hat{\Omega}(a)-\tilde{\Omega}(a)\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. Since by Theorem 2, $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$, we just need to establish HAC consitency for vectors
$\Psi_{t}=\begin{pmatrix}\hat{\psi}^{0}_{t}\\ \hat{\psi}^{1}_{t}\\ \cdots\\ \hat{\psi}^{H_{1}}_{t}\\ \hat{\psi}^{H_{1}+1}_{t}\end{pmatrix}=\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}$ (101)
The last component, $\hat{\xi}^{H_{1}+1}_{t}$, has a different structure from the other components since it is multiplied by $(Z_{t}-\bar{Z}_{t})$. From the proof of Theorem 3, we know that $\max_{t=1,\dots T}|Z_{t}-\bar{Z}_{t}|=O_{p}(T^{1/4})$. So, the conistency of the HAC components, including the linear combinations of $\hat{\psi}^{H_{1}+1}_{t}$ can be proven in a similar fashion as the other components. So, to save space, we only consider vector $a=vec(a_{0},a_{1},\dots,a_{H_{1}},0_{n})$ that puts 0 on the last $n$ components and all $a_{i}\in R^{n^{2}}$.
@@
$+\sum_{h,s=0}^{H}(\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})+\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})$ (103)
@@
The r.h.s. is $O_{p}(\sqrt{T})$ by proof of Step 2 of Lemma B.1 and Markov inequality implied by existence of $6+\epsilon$ moments of the stationary mixing sequence $\eta_{t}$.
The sums of remaining two terms in (103) are also $O_{p}(\sqrt{T})$, which can be shown using the same techniques.

### v2

@@ A.2 Smoothed LP
Figure 10: Identification of the monetary policy shock in the U.S. Smoothed IRF estimates with point-wise and simultaneous bands. Heteroscedasticity-IV smoothed LP estimator with $k=0,p=12$ (blue line) with corresponding 68% point-wise (blue dots) and simultaneous sup-t (dash line).
@@ A.3 Bonferroni bounds
Figure 11: Identification of the monetary policy shock in the U.S. Sup-t vs Bonferroni simultaneous bands. Heteroscedasticity-IV LP with $k=0,p=12$ estimator (blue solid line), Bonferroni bounds (gold dots) and simultaneous sup-t bounds (dashed line).
@@ A.4 Comparison with conventional VAR approch
Figure 12: Identification of the monetary policy shock in the U.S. LP vs VAR using Christiano et al. (1999). Cholesky identification with LP estimator with $k=0,p=12$ (blue line) and homoscedastic VAR estimator (violet solid) with corresponding 68% point-wise bands ( gold dashes and violet dots).
@@ B.1 Proof of Theorem 1
$b_{\cdot 1}^{\prime}\Sigma^{-1}_{\eta}b_{\cdot 1}=1.$ (48)
Step 2. The reduced form shocks have the following representation for any $k=1,\dots,n,$
$\eta_{1t}\eta_{kt}$ $=b_{k1}b_{11}\varepsilon^{2}_{1t}+\sum_{i\neq j}b_{ki}b_{1j}\varepsilon_{it}\varepsilon_{jt}+\sum_{i\neq 1}b_{ki}b_{1i}\varepsilon^{2}_{it}.$ (49)
@@
$=b_{\cdot 1}b_{11}\mathbb{E}(\varepsilon^{2}_{1t}-1)(Z_{t}-\mathbb{E}Z_{t})=b_{\cdot 1}b_{11}\rho_{1}.$ (52)
@@
$\sqrt{\gamma^{\prime}\Sigma^{-1}_{\eta}\gamma}=|b_{11}\rho_{1}|,$ (53)
@@
$b_{\cdot 1}$ $=\frac{1}{\sqrt{\gamma^{\prime}\Sigma_{\eta}^{-1}\gamma}}\gamma\operatorname{sign}(b_{11}\rho_{1}).$ (54)
@@ B.2 Proof of LP consistency
Suppose that Assumptions 2 and 3 hold. Consider LS residuals $\hat{\eta}_{t}(H)$ from $H-$step forward autoregression,
@@
The proof proceeds in 3 steps. First, we show that (12) can be decomposed into linearly independent univariate AR processes using Jordan decomposition. Then, we show that the LS estimators for individual processes have well-defined asymptotic distributions. Finally, we use the invariance of LS estimates with respect to linear transformations to estimate the difference in (56).
Step 1. Notice that $\mu_{t}=(1,t,\dots,t^{k})^{\prime}$ satisfies a first-order deterministic difference equation
$\mu_{t}=Q_{\mu}\mu_{t-1},$ (57)
@@
Matrix $Q_{y}$ can be represented in Jordan normal form
$Q_{y}=P_{y}J_{y}P_{y}^{-1},$ (63)
where $P_{y}$ is an invertible matrix and $J_{y}$ is a block diagonal matrix. [^22] See, for example, Gantmakher (2000). Then, following ideas in Tsay and Tiao (1990) and Sims et al. (1990), we get VAR(1) representation
@@
with $\lambda_{i}$ being $i$-th eigenvalue of $Q_{y}$ (not necessarily distinct) with geometric multiplicity $n_{i}$. Suppose, without loss of generality, that the blocks $J_{y,i}$ are arranged in decreasing order according to the absolute values of the corresponding $\lambda_{i}$. If a unit root is present, then start with $\lambda_{1}=1$. By Assumption 3, there are no roots larger than 1 in absolute value.
@@
$\tilde{X}_{t}\triangleq\begin{pmatrix}P_{x}P_{y}^{-1}Y^{\circ}_{t}\\ \mu_{t}\end{pmatrix}\triangleq F(L)\nu_{t},$ (66)
@@
and $\nu_{t}=(\eta^{\prime}_{t},(\xi^{1}_{t}(\lambda_{1}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{1}))^{\prime},(\xi^{1}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{g_{1}}_{t}(\lambda_{2}))^{\prime},\dots,(\xi^{1}_{t}(\lambda_{b}))^{\prime},\dots,(\xi^{g_{b}}_{t}(\lambda_{1}))^{\prime},1,t,\dots,t^{k})^{\prime}$ with $\xi^{j}_{t}(\lambda)\triangleq\lambda\sum_{s=1}^{t}\xi^{j-1}_{s}(\lambda)$, $\xi^{0}_{t}(\lambda)\triangleq\eta_{t}\lambda^{t}$, $i=1,\dots,b$ $\lambda_{i}$ are unit roots of (12) and $g_{i}$ are their corresponding maximum geometric multiplicities. The blocks $F_{i,j}$ corresponding to two different unit-roots $\lambda_{i}$ and $\lambda_{j}$ are 0. The lag matrix $F_{11}(L)$ is obtained by inversion of the stationary rows of (64) and therefore $\sum^{\infty}_{j=0}j^{g}|F_{11j}|<\infty$ and $\sum_{j=0}^{\infty}F_{11j}F_{11j}^{\prime}$ is not singular.
Following Sims et al. (1990) and Tsay and Tiao (1990), one can arrange the properly chosen powers of $T^{1/2}$ in a diagonal matrix $\Upsilon_{T}$ such that $\sum_{t=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{t-1}$ have a bounded variance (for example, for stationary components it is $T^{1/2}$, for components containing $\xi^{j}_{t}(\lambda_{2}))$ it is $T^{(1+j)/2}$ etc.).
@@
Consider $h=0$ first. The proof for deterministic, stationary, and ordinary unit root components follows the proofs of Lemmas 1 and 2 in Sims et al. (1990) with a modification required to account for conditionally heteroscedastic shocks $\eta_{t}$. Namely, proof of Lemma 1 Sims et al. (1990) requires replacement of Functional CLT for sums of $\xi^{1}_{t}$ in part (a) and Theorem 2.4(ii) from Chan and Wei (1988) in part (c) for case $p=1$. Under Assumption 2, Theorem 3.1 (De Jong and Davidson, 2000) applied to $\xi^{1}_{t}$ gives the required FCLT under the mixing processes assumption. Theorem 4.1 of De Jong and Davidson (2000) for any pair of components of $\xi^{1}_{t}$ replaces Theorem 2.4(ii) of Chan and Wei (1988) for the mixing sequence case (the MDS assumption on $\eta_{t}$ ensures that we do not need to recenter the sums). For complex unit roots, the limiting distribution is given in Theorem 4.4 of Tsay and Tiao (1990), which also requires substituting FCLT for martingale differences with mixing white noise (De Jong and Davidson, 2000, Theorem 4.1) and strong law of large numbers used in from Chan and Wei (1988) for MDS with Corollary 3.2 in Rio et al. (2017). For our purposes, we do not need the explicit form of $\phi$ and $\tilde{V}$.
@@
where
$\Theta(h)\triangleq(A_{1}(h),A_{2}(h),\dots,A_{p}(h),V(h))$
and
$X_{t-1}\triangleq vec(y_{t-1},y_{t-2},\dots,y_{t-p},y_{t-p-1},1,(t-1),\dots(t-1)^{k}).$
@@
$\hat{\eta}_{t+h}(h)-\eta_{t+h}(h)=(\Theta(h)-\hat{\Theta}(h))X_{t-1}=-X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h).$ (71)
@@
$X_{t-1}^{\prime}(\sum_{r=p}^{T-h}X_{r-1}X_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}X_{r-1}\eta_{r+h}(h)$ (72)
$=$ $\tilde{X}_{t-1}^{\prime}(P^{-1})^{\prime}(P^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}(P^{-1})^{\prime})^{-1}\sum_{r=p}^{T-h}(P^{-1})\tilde{X}_{r-1}\eta_{r+h}(h)$ (73)
$=$ $\tilde{X}_{t-1}^{\prime}(\sum_{r=p}^{T-h}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime})^{-1}\sum_{r=p}^{T-h}\tilde{X}_{r-1}\eta_{r+h}(h).$ (74)
@@
$=C_{h}^{\prime}+\big{(}\frac{1}{T}\sum_{t=p}^{T-h}\eta_{t}\eta_{t}^{\prime}+O_{p}(T^{-1})\big{)}^{-1}(\frac{1}{T}\sum_{t=p}^{T-h}\sum_{s=0}^{h-1}\eta_{t}(C_{s}\eta_{t+h-s})^{\prime}+O_{p}(T^{-1}))+O_{p}(T^{-1}).$ (83)
Since $\eta_{t}$ is stationary and mixing and thus sequence $\eta_{t}\eta_{t+h}$ is ergodic with finite fourth moments, we can apply the ergodic theorem to get the desired result.
@@ B.3 Proof of consistency of IV estimtor
$\bar{Z}_{T}\sum_{t=p}^{T}(\hat{\eta}_{it}-\eta_{it})\eta_{jt}=O_{p}(1).$ (86)
Next, by Assumption 2.5, $Z_{t}\eta_{t}$ is zero mean and is uncorrelated with $\tilde{X}_{t-1}^{\prime}$. Using the same argument as in Step 2 of the proof of Lemma B.1, $\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}$ is a tight random sequence by De Jong and Davidson (2000, Theorem 4.1). So
$(\sum_{t=p}^{T}(\hat{\eta}_{t}-\eta_{t})\eta^{\prime}_{t}Z_{t})^{\prime}=\sum_{t=p}^{T}Z_{t}\eta_{t}\tilde{X}_{t-1}^{\prime}\Upsilon_{T}^{-1}(\Upsilon_{T}^{-1}\sum_{r=p}^{T}\tilde{X}_{r-1}\tilde{X}_{r-1}^{\prime}\Upsilon_{T}^{-1})^{-1}\sum_{r=p}^{T}\Upsilon_{T}^{-1}\tilde{X}_{r-1}\eta^{\prime}_{r}=O_{p}(1).$ (87)
The last term in (85) can be bounded by the Cauchy-Schwarz theorem,
@@
$\sum_{t=p}^{T}\eta_{t}\eta_{1t}(Z_{t}-\bar{Z}_{T})=$ $\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})$ (93)
$+\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(Z_{t}-\mathbb{E}Z_{t})\frac{1}{\sqrt{T-p}}\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})$ (94)
$=$ $\sum_{t=p}^{T}(\eta_{t}\eta_{1t}-\mathbb{E}\eta_{t}\eta_{1t})(Z_{t}-\mathbb{E}Z_{t})+O_{p}(1).$ (95)
@@ B.4 Proof of asymptotic normality of the LP estimators
By Assumption 2, $\Xi_{t}(H)$ is a strong mixing process of size at least $-2(1+\epsilon)/\epsilon$ with zero mean and finite $2+\epsilon$ moments (condition DMR is satisfied, see p. 12 in Rio et al. (2017)). By Theorem 4.2 in Rio et al. (2017) and the Cramer-Wold theorem, there exists the long run variance matrix $\Omega$ for $\Xi_{t}(H)$ and
@@ B.5 Proof of consistency of HAC standard errors
Step 1.To show consistency of the whole matrix $\hat{\Omega}=\hat{\Omega}(H_{1})$, we need to establish consistency of HAC estimators for all combinations $a^{\prime}\Xi_{t}$. Then the HAC estimator for particular $a$ is defined as
$\hat{\Omega}(a)\triangleq\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\hat{\Xi}_{t})(\hat{\Xi}_{t+\ell}^{\prime}a).$ (101)
@@
$\tilde{\Omega}(a)\triangleq\sum_{\ell=-B_{T}}^{B_{T}}\frac{k(\ell/B_{T})}{T-\ell}\sum_{t=p}^{T-\ell}(a^{\prime}\Xi_{t})(\Xi_{t+\ell}^{\prime}a).$ (102)
@@
Step 2. What remains to show is $\hat{\Omega}(a)-\tilde{\Omega}(a)\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. Since by Theorem 2, $\hat{\Sigma}_{\eta}\stackrel{{\scriptstyle p}}{{\rightarrow}}\Sigma_{\eta}$, we just need to establish HAC consistency for vectors
$\Psi_{t}=\begin{pmatrix}\hat{\psi}^{0}_{t}\\ \hat{\psi}^{1}_{t}\\ \cdots\\ \hat{\psi}^{H_{1}}_{t}\\ \hat{\psi}^{H_{1}+1}_{t}\end{pmatrix}=\begin{pmatrix}vec(\hat{\eta}_{t}\hat{\eta}_{t}^{\prime}-\hat{\Sigma}_{\eta})\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+1}(1)}))\\ \cdots\\ vec((\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)-\overline{\hat{\eta}_{t}\hat{\eta}^{\prime}_{t+h}(H)}))\\ (\hat{\eta}_{t}\hat{\eta}_{1t}-\overline{\hat{\eta}_{t}\hat{\eta}_{1t}})(Z_{t}-\bar{Z}_{t})-\hat{\gamma}\end{pmatrix}.$ (103)
The last component, $\hat{\xi}^{H_{1}+1}_{t}$, has a different structure from the other components since it is multiplied by $(Z_{t}-\bar{Z}_{t})$. From the proof of Theorem 3, we know that $\max_{t=1,\dots T}|Z_{t}-\bar{Z}_{t}|=O_{p}(T^{1/4})$. So, the consistency of the HAC components, including the linear combinations of $\hat{\psi}^{H_{1}+1}_{t}$ can be proven similarly to the other components. So, to save space, we only consider vector $a=vec(a_{0},a_{1},\dots,a_{H_{1}},0_{n})$ that puts 0 on the last $n$ components and all $a_{i}\in R^{n^{2}}$.
@@
$+\sum_{h,s=0}^{H}(\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell})$ (105)
$+\sum_{h,s=0}^{H}(\hat{\psi}^{h}_{t}-\psi^{h}_{t})^{\prime}a_{h}a^{\prime}_{s}(\hat{\psi}^{s}_{t+\ell}-\psi^{s}_{t+\ell}).$ (106)
@@
The r.h.s. is $O_{p}(\sqrt{T})$ by proof of Step 2 of Lemma B.1 and Markov inequality implied by the existence of $6+\epsilon$ moments of the stationary mixing sequence $\eta_{t}$.
The sums of the remaining two terms in (106) are also $O_{p}(\sqrt{T})$, which can be shown using the same techniques.